A block of gold has length 5.62 cm, width 6.35 cm, and height 2.78 cm. . Now multiply the rounded result of part (a) by the height and again round, obtaining the volume. Let this be 𝑉1. 


1
Expert's answer
2021-01-26T08:41:10-0500

The area of the block's base is:


"A = L\times W"

where "L = 5.62cm" is the length, and "W = 6.35cm" is the width. Thus, obtain:


"A = 35.687cm^2\approx35.69cm^2"

The volume is then:


"V1= A\times H"

where "H = 2.78cm" is the height. Thus, obtain:


"V1 = 35.69cm^2\times 2.78cm =99.2182 cm^3 \approx 99.22cm^3"

Answer. 99.22 cm^3.